# Optimal Control of Behavioral-Feedback SIR Epidemic Model

### v1

@@ 1 Introduction
In recent years, mathematical models have become important tools for predicting how outbreaks evolve and for assessing the effectiveness of different control strategies. Deterministic compartmental models play a central role in capturing the essential features of disease transmission and supporting the formulation of public health strategies. Among these, the Susceptible-Infected-Recovered (SIR) model is one of the most extensively studied because of its simplicity and generality [Kermack.McKendrick:1927, Hethcote2000TheMO, MKendrickApplicationsOM, Diekmann2000MathematicalEO]. It divides the population into three compartments (susceptible, infected, and recovered) and describes their evolution in time through a system of nonlinear differential equations, assuming that the population is large, homogeneous, and well mixed.
@@
However, a key limitation of the classical SIR model is its assumption of a constant infection rate, disregarding behavioral adaptations. Historical and recent evidence shows that agents change their behavior in response to perceived infection risk, affecting contact patterns and, consequently, the disease transmission. For example, during the 1918 influenza pandemic people avoided crowded areas [crosby2003america], during the SARS outbreak, widespread use of face masks and reduced travel were observed [lau2005sars], and during the recent COVID-19 pandemic behavioral responses were further reinforced by government measures such as social distancing and lockdowns [hsiang2020effect].
@@
To incorporate the human behavior into epidemic models, two main approaches have emerged. The first one consists in adding to the system new dynamic variables, often based on evolutionary game theory [sandholm2010population], to model how the agents weigh the costs and benefits of actions like vaccination or isolation based on their risk perception and on social influence. The infection rate is then assumed to depend on such variables, which in turn depend on the epidemic state [satapathi2022epidemic, ye2021game, frieswijk2022meanfield, paarporn2023sis, amaral2021epidemiological, kabir2020evolutionary, martins2023epidemic, certorio2022epidemic]. The drawback of this approach is that the number of variables of the system increases, making optimal control problems often analytically intractable.
@@
The second main modeling approach incorporates the human behavior through static feedback mechanisms. In these behavioral-feedback SIR epidemic models, the infection rate $\beta$ is a function of the current epidemic state rather than dependent on additional dynamic variables or constant as in the classical SIR model  [funk2010modelling, verelst2016review]. A seminal contribution for this class of models is [CAPASSO197843], which introduced a SIR model where the infection rate $\beta$ decreases with the fraction of infected agents $y$, producing a unimodal infection curve. Later studies confirmed that such feedback can flatten the curve and preserve key stability properties [Baker2020, franco2020feedback, korobeinikov2006lyapunov], and that the unimodality of the infection curve is preserved even if the the infection rate depends on both the fraction of susceptible agents $x$ and infected agents $y$ [Alutto2021OnSE]. Recent work further explored these effects, deriving bounds on epidemic overshoot and identifying invariants of motion [nguyen2024upper], or studying complex nonlinear feedback with saturation effects [srivastava2024nonlinear]. Other behavioral-feedback SIR models with non-linear infection rate $\beta(x,y)$ are considered in [feng2000endemic, gao2024final, liu1987dynamical]. Time-dependent infection rates have also been introduced to reflect seasonal or delayed responses [boatto2018sir].
@@
In this paper, we address an optimal control problem for a class of behavioral-feedback SIR models where the infection rate depends on both the fraction of susceptible agents $x$ and infected agents $y$. Optimal control of the classical SIR epidemic model has been extensively studied in recent years [zino2021analysis, yi2022edge, morton1974optimal, behncke2000optimal, parino2024optimal], particularly for calibrating lockdowns to achieve epidemic control without excessively harming the economy. For example, [kruse2020optimal] proposes an optimal control problem for the SIR model over a finite time horizon, where the epidemic cost is linear in the fraction of infected agents, and the economic cost is linear in the level of intervention, proving that the optimal control is bang-bang with at most two switches. In [cianfanelli2021lockdown], the authors study an optimal control problem where the economic cost depends on both the severity and duration of the lockdown, and the epidemic cost is quadratic in the fraction of infected agents to account for hospital congestion. They prove that, for an infinite time horizon, a strategy that stabilizes the fraction of infected agents outperforms one that stabilizes the reproduction number. Similarly, in [acemoglu2021optimal], the authors solve an optimal control problem numerically for network SIR dynamics, showing that optimal policies differentially targeting risk or age groups significantly outperform uniform policies. In [Miclo.ea:2022] the authors provide the solution of an optimal control problem of the SIR model with infinite time horizon, where an external controller aims to minimize the cost of the intervention while ensuring that the fraction of infected agents remains below a specified threshold at every time. This threshold represents the healthcare system’s capacity, referred to as the ICU capacity constraint, ensuring comprehensive care for infected patients. Such constraints reflect scenarios like those observed during COVID-19, where the number of critically ill patients risked exceeding available resources, potentially leading to adverse outcomes. Respecting the ICU constraint guarantees that all patients receive necessary medical care. The optimal strategy consists in setting the minimum lockdown level needed to satisfy the constraint, known as the filling the box strategy. In the first stages, the disease spreads uncontrolled, followed by an abrupt lockdown when the infection curves meets the threshold, and then a gradual reopening until a time after which the spread is no longer regulated and the infection naturally fades out. The economic rationale behind this optimal policy is to avoid shutting down society while ICU resources remain unused. Therefore, whenever the natural spread does not threaten the ICU constraint, it should proceed without intervention. This policy is proven to be optimal also in [Acemoglu2021OptimalAT], where the authors propose an optimal control problem for epidemic mitigation that incorporates both molecular and serology testing.
@@
Our main contribution is to prove that, under some mild assumptions on the regularity and monotonicity of feedback mechanism that regulates the infection rate, the filling-the-box strategy is the optimal control for a general class of behavioral-feedback SIR models. This result extends the findings of [Miclo.ea:2022], which apply to the standard SIR model. The proof relies on the properties of the behavioral-feedback SIR model, in particular the unimodality of the infection curve, which is guaranteed under very general assumptions on the feedback mechanism. Compared to [Miclo.ea:2022], the main technical challenge of our analysis is that the class of behavioral feedback SIR models considered in this paper does not admit in general an invariant of motion, which was exploited in [Miclo.ea:2022] to prove the optimality of the filling-the-box strategy.
To the best of our knowledge, this is the first study that jointly addresses the human adaptive behavior and the optimal control of epidemics besides our preliminary results in [CDC2025], where the authors study a behavioral-feedback SIR model with a special feedback mechanism, showing that it admits a non-trivial invariant of motion and characterize the infection peak. In addition, [CDC2025] analyzes the system under a threshold constraint on the infection level, showing that a feasible control strategy consists in applying the minimal amount of restrictions required to keep the epidemic below a threshold, and they provide an explicit characterization of the associated cost. The main novelties of this work with respect to [CDC2025] are the following. First, we consider a general class of feedback mechanisms, rather than focusing on a specific functional form. Within this generalized framework, we prove that under mild conditions on the transmission rate the infection curve remains unimodal, and we rigorously formulate the associated optimal control problem. In this setting, we establish that the filling-the-box strategy is the optimal solution, thus extending the preliminary results of [CDC2025] to a much broader class of models, where the analysis cannot rely on invariants of motion and requires different arguments. Finally, we provide a numerical counterexample showing that, when the monotonicity assumptions on the infection rate are relaxed, the filling-the-box strategy is no longer optimal.
@@
The rest of the paper is organized as follows. In Section 2 we introduce the controlled behavioral-feedback SIR epidemic model, formulate the optimal control problem we aim to study, and present our main result. In Section 3, we analyze the uncontrolled dynamics, which is instrumental for the proof of our main result. The proof itself is then provided in Section 4. In Section 5, we discuss a counterexample, while in Section 6 we summarize our contribution and discuss future research lines.
@@ 1.1 Notation
We briefly gather here some notational conventions adopted throughout the paper. We denote by $\mathbb{R}$ and $\mathbb{R}_{+}$ the sets of real and nonnegative real numbers, respectively. We will often consider differentiable functions $f(x,y)$ and use the notation $\dot{f}=f_{x}\dot{x}+f_{y}\dot{y}$ to denote the time derivative of the composed function $f(x(t),y(t))$. Throughout this paper, given an interval $\mathcal{I}\subseteq\mathbb{R}$, we say that $u:\mathcal{I}\to\mathbb{R}^{n}$ is a piecewise-continuous function if there exists a subset $\mathcal{W}\subseteq\mathcal{I}$ that does not contain accumulation points and is such that:
@@
A function $u:\mathcal{I}\to\mathbb{R}^{n}$ is a right-continuous if it is piecewise-continuous and $u(t_{0})=u(t_{0}^{+})$ for every discontinuity point $t_{0}$. Furthermore, we say that a continuous function $u:\mathcal{I}\to\mathbb{R}$ is piecewise-$\mathcal{C}^{1}$ if, except for a subset $\mathcal{W}$ as above, it is differentiable and $u^{\prime}$ is piecewise-continuous.
@@ 2.1 Model Definition
    $t\to x(t)$ is non-increasing for every $t\geq 0$ and is strictly decreasing for every $t\geq 0$ if and only if $x_{0}>0$ and $y_{0}>0$;
@@
From now on, we will refer to the vector functions constructed in the previous proposition as to solutions of system (1).
@@
Equation. (5) captures how the infection rate varies with the fraction of susceptible agents. This condition is satisfied when $\beta_{x}(x,y)$ is positive, e.g., when a higher fraction of susceptible leads to a lower perceived risk, encouraging social interactions and riskier behaviors that facilitate the transmission of the disease. Equation (5) can also be satisfied when $\beta_{x}(x,y)$ is negative but not too large in magnitude. A negative value of $\beta_{x}(x,y)$ can be interpreted as the effect of pandemic fatigue, where agents gradually relax self-protective measures as the epidemic progresses. Indeed, by Proposition 1(i), the fraction of susceptible agents $x$ decreases over time, so this condition reflects an increase in the infection rate as time progresses. Equation (6) captures the tendency of the agents to reduce their interaction level in response to the number of active cases. We remark that (5) generalizes the assumption made in [Alutto2021OnSE], while (6) is a standard assumption in the literature of feedback SIR epidemic models [Baker2020, franco2020feedback].
@@
satisfy (5)-(6). The CBF-SIR epidemic model with infection rate (7) has been studied in [CDC2025]. Note that also the classical SIR model may be recovered as a special case from (7)-(8) by letting $a=0$ and $b$ being a positive constant.
@@ 2.2 Problem Statement
Note that the solution of (9) is trivial if $y_{0}=0$, since Proposition 1(ii) guarantees that in this case $y(t)=0$ for every $t\geq 0$ under every control signal in $\mathcal{U}$, including the zero cost control $u\equiv 0$. For this reason, we will work under the assumption that $y_{0}>0$. In the next section we establish our main result, which provides a solution of (9).
@@ 2.3 Main result
hence the fraction of infected remains constantly at the threshold $\bar{y}$. At time $T_{1}$ such that $R(x^{*}(T_{1}),\bar{y})=1$ (or equivalently, $\rho(x^{*}(T_{1}),\bar{y})=0$), the control is released, so that the trajectory follows the uncontrolled dynamics and the infection curve naturally declines from there on. The corresponding control signal $u^{*}(t)$ starts abruptly above 0 at time $T_{0}$ and subsequently is adjusted in time to maintain the fraction of infected exactly at the threshold $\bar{y}$ until $T_{1}$. At that point, the control is deactivated. This policy is known in the literature as filling-the-box strategy and its optimality relies on the monotonicity of the reproduction number, which implies the unimodality of the infection curve. Our result generalizes [Miclo.ea:2022], which proved that the filling-the-box strategy is the optimal control of the equivalent optimal control problem formulated for the classical SIR epidemic model, which is a special case of the CBF-SIR epidemic model, as noted in Example 2.3.
@@
The next section analyzes the property of the CBF-SIR epidemic model without control. This analysis will be instrumental for the proof of Theorem 2.5.
@@ 3.1 Unimodality of the infection curve
We start by proving that the infection curve of the uncontrolled BF-SIR epidemic model (14) under condition (5) is unimodal, namely, it admits at most one peak. This result generalizes [Alutto2021OnSE, Theorem 1].
@@
Note that, for a given time $\hat{t}$ in which $R(x(\hat{t}),y(\hat{t}))=1$, this equation, together with (5), implies that $\dot{R}(x(\hat{t}),y(\hat{t}))<0$. This proves that all stationary points of $y(t)$ are local maxima. Hence, if $R(x_{0},y_{0})\leq 1$, it must be $R(x(t),y(t))<1$ for every $t>0$, so that $y(t)$ is necessarily strictly decreasing for every positive time $t\geq 0$, proving item (i). To prove item (ii), let $R(x_{0},y_{0})>1$ and assume by contradiction that $R(x(t),y(t))>1$ for every $t\geq 0$. Then, $y(t)$ would be strictly increasing for all times and could not converge to an equilibrium point, contradicting Proposition 1(iv). Therefore, $\hat{t}=\min\{t\geq 0\,|\,R(x(t),y(t))=1\}$ must exist and has to be finite, otherwise the fraction of recovered agents would diverge in time due to the third equation in (1), violating Proposition 1. Hence, $y(t)$ is strictly increasing in $0,\hat{t}$]. For $t\geq\hat{t}$, item (i) implies that $y(t)$ is strictly decreasing, proving item (ii).
@@ 3.2 Some geometric considerations
It follows from [thesisMartina, Chapter 2] that, for every initial condition $(x_{0},y_{0})$ in $\mathcal{S}$, the uncontrolled BF-SIR model (14) admits a unique solution of class $\mathcal{C}^{1}$ in $\mathcal{S}$. Consequently, there exists a map $\phi:\mathbb{R}_{+}\times\mathcal{S}\to\mathcal{S}$ such that $(x_{0},y_{0})\mapsto\phi(t,x_{0},y_{0})$ and $t\mapsto\phi(t,x_{0},y_{0})$ are both of class $\mathcal{C}^{1}$ and $t\mapsto\phi(t,x_{0},y_{0})$ is the unique solution of the uncontrolled BF-SIR model (14) with $\phi(0;x_{0},y_{0})=(x_{0},y_{0})$ in $\mathcal{S}$. The positive semi-orbit for the initial condition $(x_{0},y_{0})$ is then defined as
@@
The first part of our analysis leads to a decomposition of $\mathcal{D}_{\bar{y}}$ into two sets $\mathcal{D}_{\bar{y}}^{-}$ and $\mathcal{D}_{\bar{y}}^{+}$, which will be formally defined after Lemma 3.12. Informally, $\mathcal{D}_{\bar{y}}^{-}$ contains the initial conditions from which the fraction of infected remains below the threshold $\bar{y}$ without requiring an external intervention and is the largest positively invariant set contained in $\mathcal{D}_{\bar{y}}$, as established by Lemma 3.16. In contrast, $\mathcal{D}_{\bar{y}}^{+}$ contains all initial conditions such that the solution of the uncontrolled dynamics would exceed the threshold $\bar{y}$, thus requiring an external intervention. The two regions are illustrated in Figure 2.
@@
Note from (3) that $R(0,1)=0$, while $R(1,0)>1$ due to (16). Therefore, since $y\mapsto R(1-y,y)$ is continuous on $[0,1]$, there exists $\tilde{y}$ in $(0,1)$ such that $R(1-\tilde{y},\tilde{y})=1$. For every $y\in[0,\tilde{y}$], note that $R(0,y)=0$ by (3). Moreover, for every such $y$, since $R(1-\tilde{y},\tilde{y})=1$, it follows from (5)-(6) that $R(1-y,y)\geq 1$. Therefore, (5) implies the existence of a unique $x=\kappa(y)$ such that $R(\kappa(y),y)=1$. The fact that $\kappa$ is of class $\mathcal{C}^{1}$ follows from the Implicit Function Theorem [rudin1976principles, Theorem 9.18] and from (5).
@@
    $\lambda(y)>\kappa(y)$ for every $y$ in $\hat{y},\bar{y})$ and $\lambda(\bar{y})=\kappa(\bar{y})=\bar{x}$. In particular, $R(\lambda(y),y)>1$ for every $\hat{y},\bar{y})$;
@@
We define the dynamics
$\begin{cases}\dot{x}=\beta(x,y)xy,\\ \dot{y}=-\beta(x,y)xy+\gamma y\,.\end{cases}$ (17)
@@
The negative semi-orbit $\Gamma^{-}(\bar{x},\bar{y})$ is represented by the thick blue curve in Figure 2. The green and the light-blue area are, respectively, the sets $\mathcal{D}_{\bar{y}}^{-}$ and $\mathcal{D}_{\bar{y}}^{+}$. If the negative semi-orbit that starts from $(\bar{x},\bar{y})$ intercepts a point $(x,y)$ on the upper-right boundary of $\mathcal{S}$, then $\hat{y}>0$, otherwise $\hat{y}=0$.
@@
Let Assumption 1 hold true. Then, if (16) does not hold true, $\tilde{y}$ and $\hat{y}$ are not defined and $R(x,y)\leq 1$ for every $(x,y)$ in $\mathcal{S}$. Hence, the set $\mathcal{D}_{\bar{y}}^{+}$ is empty by Remark 3.14 and $\mathcal{D}_{\bar{y}}=\mathcal{D}_{\bar{y}}^{-}$. Furthermore, the set $\mathcal{D}_{\bar{y}}^{+}$ is empty if and only if $R(1-\bar{y},\bar{y})\leq 1$. If (16) holds true, this is equivalent to $\tilde{y}\leq\bar{y}$. If in particular $\tilde{y}<\bar{y}$, then $\bar{x}$ does not exist, therefore neither $\hat{y}$ and the function $\lambda$ are well defined. If instead $\tilde{y}=\bar{y}$, then $\bar{x}=1-\bar{y}$, $\hat{y}=\tilde{y}=\bar{y}$ and the curve $x=\lambda(y)$ degenerates to a point $(1-\bar{y},\bar{y})$.

### v2

@@ 1 Introduction
In recent years, mathematical models have become important tools for predicting how outbreaks evolve and for assessing the effectiveness of different control strategies. Deterministic compartmental models play a central role in capturing the essential features of disease transmission and supporting the formulation of public health strategies. Among these, the Susceptible-Infected-Recovered (SIR) model is one of the most extensively studied because of its simplicity and generality [26, 22, 34, 14]. It divides the population into three compartments (susceptible, infected, and recovered) and describes their evolution in time through a system of nonlinear differential equations, assuming that the population is large, homogeneous, and well mixed.
@@
However, a key limitation of the classical SIR model is its assumption of a constant infection rate, disregarding behavioral adaptations. Historical and recent evidence shows that agents change their behavior in response to perceived infection risk, affecting contact patterns and, consequently, the disease transmission. For example, during the 1918 influenza pandemic people avoided crowded areas [13], during the SARS outbreak, widespread use of face masks and reduced travel were observed [30], and during the recent COVID-19 pandemic behavioral responses were further reinforced by government measures such as social distancing and lockdowns [23].
@@
To incorporate the human behavior into epidemic models, two main approaches have emerged. The first one consists in adding to the system new dynamic variables, often based on evolutionary game theory [41], to model how the agents weigh the costs and benefits of actions like vaccination or isolation based on their risk perception and on social influence. The infection rate is then assumed to depend on such variables, which in turn depend on the epidemic state [42, 45, 17, 38, 6, 24, 33, 11]. The drawback of this approach is that the number of variables of the system increases, making optimal control problems often analytically intractable.
@@
The second main modeling approach incorporates the human behavior through static feedback mechanisms. In these behavioral-feedback SIR epidemic models, the infection rate $\beta$ is a function of the current epidemic state rather than dependent on additional dynamic variables or constant as in the classical SIR model  [18, 44]. A seminal contribution for this class of models is [10], which introduced a SIR model where the infection rate $\beta$ decreases with the fraction of infected agents $y$, producing a unimodal infection curve. Later studies confirmed that such feedback can flatten the curve and preserve key stability properties [7, 16, 28], and that the unimodality of the infection curve is preserved even if the the infection rate depends on both the fraction of susceptible agents $x$ and infected agents $y$ [4]. Recent work further explored these effects, deriving bounds on epidemic overshoot and identifying invariants of motion [37], or studying complex nonlinear feedback with saturation effects [43]. Other behavioral-feedback SIR models with non-linear infection rate $\beta(x,y)$ are considered in [15, 19, 32]. Time-dependent infection rates have also been introduced to reflect seasonal or delayed responses [9].
@@
In this paper, we address an optimal control problem for a class of behavioral-feedback SIR models where the infection rate depends on both the fraction of susceptible agents $x$ and infected agents $y$. Optimal control of the classical SIR epidemic model has been extensively studied in recent years [47, 46, 36, 8, 39], particularly for calibrating lockdowns to achieve epidemic control without excessively harming the economy. For example, [29] proposes an optimal control problem for the SIR model over a finite time horizon, where the epidemic cost is linear in the fraction of infected agents, and the economic cost is linear in the level of intervention, proving that the optimal control is bang-bang with at most two switches. In [12], the authors study an optimal control problem where the economic cost depends on both the severity and duration of the lockdown, and the epidemic cost is quadratic in the fraction of infected agents to account for hospital congestion. They prove that, for an infinite time horizon, a strategy that stabilizes the fraction of infected agents outperforms one that stabilizes the reproduction number. Similarly, in [1], the authors solve an optimal control problem numerically for network SIR dynamics, showing that optimal policies differentially targeting risk or age groups significantly outperform uniform policies. In [35] the authors provide the solution of an optimal control problem of the SIR model with infinite time horizon, where an external controller aims to minimize the cost of the intervention while ensuring that the fraction of infected agents remains below a specified threshold at every time. This threshold represents the healthcare system’s capacity, referred to as the ICU capacity constraint, ensuring comprehensive care for infected patients. Such constraints reflect scenarios like those observed during COVID-19, where the number of critically ill patients risked exceeding available resources, potentially leading to adverse outcomes. Respecting the ICU constraint guarantees that all patients receive necessary medical care. The optimal strategy consists in setting the minimum lockdown level needed to satisfy the constraint, known as the filling the box strategy. In the first stages, the disease spreads uncontrolled, followed by an abrupt lockdown when the infection curves meets the threshold, and then a gradual reopening until a time after which the spread is no longer regulated and the infection naturally fades out. The economic rationale behind this optimal policy is to avoid shutting down society while ICU resources remain unused. Therefore, whenever the natural spread does not threaten the ICU constraint, it should proceed without intervention. This policy is proven to be optimal also in [2], where the authors propose an optimal control problem for epidemic mitigation that incorporates both molecular and serology testing.
@@
Our main contribution is to prove that, under some mild assumptions on the regularity and monotonicity of feedback mechanism that regulates the infection rate, the filling-the-box strategy is the optimal control for a general class of behavioral-feedback SIR models. This result extends the findings of [35], which apply to the standard SIR model. The proof relies on the properties of the behavioral-feedback SIR model, in particular the unimodality of the infection curve, which is guaranteed under very general assumptions on the feedback mechanism. Compared to [35], the main technical challenge of our analysis is that the class of behavioral feedback SIR models considered in this paper does not admit in general an invariant of motion, which was exploited in [35] to prove the optimality of the filling-the-box strategy.
This work extends our preliminary results in [3]. Therein, we i) derived a non-trivial invariant of motion, ii) characterized the infection peak and iii) computed the cost of filling-the-box strategy for a behavioral-feedback SIR model with a specific feedback mechanism. This work extends such results by considering a general class of feedback mechanisms, rather than focusing on a specific functional form and by proving optimality of the filling-the-box strategy (without the need of an invariant of motion for the analysis). Moreover, we provide a numerical counterexample showing that, when the monotonicity assumptions on the infection rate are relaxed, the filling-the-box strategy is no longer optimal, thus proving that it is not possible to further extend the optimality result.
@@
The rest of the paper is organized as follows. In Section 2 we introduce the controlled behavioral-feedback SIR epidemic model, formulate the optimal control problem we aim to study, and present our main result. In Section 3, we analyze the uncontrolled dynamics, as a first instrumental step for the proof of our main result. The proof itself is then provided in Section 4. In Section 5, we discuss a counterexample, while in Section 6 we summarize our contribution and discuss future research lines.
@@ 1.1 Notation
We briefly gather here some notational conventions adopted throughout the paper. We denote by $\mathbb{R}$ and $\mathbb{R}_{+}$ the sets of real and nonnegative real numbers, respectively. We will often consider differentiable functions $f(x,y)$ and use the notation $\dot{f}=f_{x}\dot{x}+f_{y}\dot{y}$ to denote the time derivative of the composite function $f(x(t),y(t))$. Throughout this paper, given an interval $\mathcal{I}\subseteq\mathbb{R}$, we say that $u:\mathcal{I}\to\mathbb{R}^{n}$ is a piecewise-continuous function if there exists a subset $\mathcal{W}\subseteq\mathcal{I}$ that does not contain accumulation points and is such that:
@@
A function $u:\mathcal{I}\to\mathbb{R}^{n}$ is right-continuous if it is piecewise-continuous and $u(t_{0})=u(t_{0}^{+})$ for every discontinuity point $t_{0}$. Furthermore, we say that a continuous function $u:\mathcal{I}\to\mathbb{R}$ is piecewise-$\mathcal{C}^{1}$ if, except for a subset $\mathcal{W}$ as above, it is differentiable and $u^{\prime}$ is piecewise-continuous.
@@ 2.1 Model Definition
    $t\to x(t)$ is non-increasing and is strictly decreasing if and only if $x_{0}>0$ and $y_{0}>0$;
@@
From now on, we will refer to the vector functions constructed in the previous proposition as solutions of system (1).
@@
Equation. (5) captures how the infection rate varies with the fraction of susceptible agents. This condition is satisfied for example when $\beta_{x}(x,y)$ is positive (e.g., when a higher fraction of susceptible leads to a lower perceived risk, encouraging social interactions and riskier behaviors that facilitate the transmission of the disease). Equation (5) can also be satisfied when $\beta_{x}(x,y)$ is negative but not too large in magnitude. A negative value of $\beta_{x}(x,y)$ can be interpreted as the effect of pandemic fatigue, where agents gradually relax self-protective measures as the epidemic progresses. Indeed, by Proposition 1(i), the fraction of susceptible agents $x$ decreases over time, so this condition reflects an increase in the infection rate as time progresses. Equation (6) captures the tendency of the agents to reduce their interaction level in response to the number of active cases. We remark that (5) generalizes the assumption made in [4], while (6) is a standard assumption in the literature of feedback SIR epidemic models [7, 16].
@@
satisfy (5)-(6). The CBF-SIR epidemic model with infection rate (7) has been studied in [3]. Note that also the classical SIR model may be recovered as a special case from (7)-(8) by letting $a=0$ and $b$ being a positive constant.
@@ 2.2 Problem Statement
Note that the solution of (9) is trivial if $y_{0}=0$, since Proposition 1(ii) guarantees that in this case $y(t)=0$ for every $t\geq 0$ under every control signal in $\mathcal{U}$, including the zero cost control $u\equiv 0$. For this reason, we will work under the assumption that $y_{0}>0$. In the next section we establish our main result, which provides a solution to (9).
@@ 2.3 Main result
hence the fraction of infected remains constantly at the threshold $\bar{y}$. At time $T_{1}$ such that $R(x^{*}(T_{1}),\bar{y})=1$ (or equivalently, $\rho(x^{*}(T_{1}),\bar{y})=0$), the control is released, so that the trajectory follows the uncontrolled dynamics and the infection curve naturally declines from there on. The corresponding control signal $u^{*}(t)$ starts abruptly above 0 at time $T_{0}$ and subsequently is adjusted in time to maintain the fraction of infected exactly at the threshold $\bar{y}$ until $T_{1}$. At that point, the control is deactivated. This policy is known in the literature as filling-the-box strategy and its optimality relies on the monotonicity of the reproduction number, which implies the unimodality of the infection curve. Our result generalizes [35], which proved that the filling-the-box strategy is the optimal control for the classical SIR epidemic model, which is a special case of the CBF-SIR epidemic model, as noted in Example 2.3.
@@
The next section analyzes the property of the CBF-SIR epidemic model without control. This analysis is instrumental for the proof of Theorem 2.5.
@@ 3.1 Unimodality of the infection curve
We start by proving that the infection curve of the uncontrolled BF-SIR epidemic model (14) under condition (5) is unimodal, namely, it admits at most one peak. This result generalizes [4, Theorem 1].
@@
Note that, for a given time $\hat{t}$ in which $R(x(\hat{t}),y(\hat{t}))=1$, this equation, together with (5), implies that $\dot{R}(x(\hat{t}),y(\hat{t}))<0$. Hence, if $R(x_{0},y_{0})\leq 1$, it must be $R(x(t),y(t))<1$ for every $t>0$, so that $y(t)$ is necessarily strictly decreasing for every positive time $t\geq 0$, proving item (i). To prove item (ii), let $R(x_{0},y_{0})>1$ and assume by contradiction that $R(x(t),y(t))>1$ for every $t\geq 0$. Then, $y(t)$ would be strictly increasing for all times and could not converge to zero, contradicting Proposition 1(iv). Therefore, $\hat{t}=\min\{t\geq 0\,|\,R(x(t),y(t))=1\}$ must exist and has to be finite, otherwise the fraction of recovered agents would diverge in time due to the third equation in (1), violating Proposition 1. Hence, $y(t)$ is strictly increasing in $0,\hat{t}$]. For $t\geq\hat{t}$, item (i) implies that $y(t)$ is strictly decreasing, proving item (ii).
@@ 3.2 Some geometric considerations
It follows from [5, Chapter 2] that, for every initial condition $(x_{0},y_{0})$ in $\mathcal{S}$, the uncontrolled BF-SIR model (14) admits a unique solution of class $\mathcal{C}^{1}$ in $\mathcal{S}$. Consequently, there exists a map $\phi:\mathbb{R}_{+}\times\mathcal{S}\to\mathcal{S}$ such that $(x_{0},y_{0})\mapsto\phi(t,x_{0},y_{0})$ and $t\mapsto\phi(t,x_{0},y_{0})$ are both of class $\mathcal{C}^{1}$ and $t\mapsto\phi(t,x_{0},y_{0})$ is the unique solution of the uncontrolled BF-SIR model (14) with $\phi(0,x_{0},y_{0})=(x_{0},y_{0})$ in $\mathcal{S}$. The positive semi-orbit for the initial condition $(x_{0},y_{0})$ is defined as
@@
The first part of our analysis leads to a decomposition of $\mathcal{D}_{\bar{y}}$ into two sets $\mathcal{D}_{\bar{y}}^{-}$ and $\mathcal{D}_{\bar{y}}^{+}$, which will be formally defined after Lemma 3.12. Informally, $\mathcal{D}_{\bar{y}}^{-}$ contains the initial conditions from which the fraction of infected remains below the threshold $\bar{y}$ without requiring an external intervention and is the largest positively invariant set contained in $\mathcal{D}_{\bar{y}}$, as established by Lemma 3.16. In contrast, $\mathcal{D}_{\bar{y}}^{+}=\mathcal{D}_{\bar{y}}\setminus\mathcal{D}_{\bar{y}}^{-}$ contains all initial conditions such that the solution of the uncontrolled dynamics would exceed the threshold $\bar{y}$, thus requiring an external intervention. The two regions are illustrated in Figure 2.
@@
Note from (3) that $R(0,1)=0$, while $R(1,0)>1$ due to (16). Therefore, since $y\mapsto R(1-y,y)$ is continuous on $[0,1]$, there exists $\tilde{y}$ in $(0,1)$ such that $R(1-\tilde{y},\tilde{y})=1$. For every $y\in[0,\tilde{y}$], note that $R(0,y)=0$ by (3). Moreover, for every such $y$, since $R(1-\tilde{y},\tilde{y})=1$, it follows from (5)-(6) that $R(1-y,y)\geq 1$. Therefore, (5) implies the existence of a unique $x=\kappa(y)$ such that $R(\kappa(y),y)=1$. The fact that $\kappa$ is of class $\mathcal{C}^{1}$ follows from the Implicit Function Theorem [40, Theorem 9.18] and from (5).
@@
    $\lambda(y)>\kappa(y)$ for every $y$ in $\hat{y},\bar{y})$ and $\lambda(\bar{y})=\kappa(\bar{y})=\bar{x}$. In particular, $R(\lambda(y),y)>1$ for every $\hat{y},\bar{y})$.
@@
We define the “backwards” dynamics
[…]
@@
The negative semi-orbit $\Gamma^{-}(\bar{x},\bar{y})$ is represented by the thick blue curve in Figure 2. The green and the light-blue area are, respectively, the sets $\mathcal{D}_{\bar{y}}^{-}$ and $\mathcal{D}_{\bar{y}}^{+}$. If the negative semi-orbit that starts from $(\bar{x},\bar{y})$ intercepts a point $(x,y)$ on the right boundary of $\mathcal{S}$, then $\hat{y}>0$, otherwise $\hat{y}=0$.
@@
Let Assumption 1 hold true. The analysis above holds under (16) and $\bar{y}\leq\tilde{y}$. If (16) does not hold, $\tilde{y}$ and $\hat{y}$ are not defined and $R(x,y)\leq 1$ for every $(x,y)$ in $\mathcal{S}$. Hence, the set $\mathcal{D}_{\bar{y}}^{+}$ is empty by Remark 3.14 and $\mathcal{D}_{\bar{y}}=\mathcal{D}_{\bar{y}}^{-}$. Instead if (16) holds and $\bar{y}>\tilde{y}$, then $\bar{x}$ does not exist, therefore neither $\hat{y}$, the function $\lambda$ and the sets $\mathcal{D}_{\bar{y}}^{+}$ and $\mathcal{D}_{\bar{y}}^{-}$ are well defined. In both cases, the optimal control is trivial because the uncontrolled trajectory will stay in $\mathcal{D}_{\bar{y}}$.